# SMART: Self-Aware Agent for Tool Overuse Mitigation

### v1

@@ Abstract
Current Large Language Model (LLM) agents demonstrate strong reasoning and tool use capabilities, but often lack self-awareness, failing to balance these approaches effectively. This imbalance leads to Tool Overuse, where models unnecessarily rely on external tools for tasks solvable with parametric knowledge, increasing computational overhead. Inspired by human metacognition, we introduce SMART (Strategic Model-Aware Reasoning with Tools), a paradigm that enhances an agent’s self-awareness to optimize task handling and reduce tool overuse. To support this paradigm, we introduce SMART-ER, a dataset spanning three domains, where reasoning alternates between parametric knowledge and tool-dependent steps, with each step enriched by rationales explaining when tools are necessary. Through supervised training, we develop SMARTAgent, a family of models that dynamically balance parametric knowledge and tool use. Evaluations show that SMARTAgent reduces tool use by 24% while improving performance by over 37%, enabling 7B-scale models to match its 70B counterpart and GPT-4o. Additionally, SMARTAgent generalizes to out-of-distribution test data like GSM8K and MINTQA, maintaining accuracy with just one-fifth the tool calls. These highlight the potential of strategic tool use to enhance reasoning, mitigate overuse, and bridge the gap between model size and performance, advancing intelligent and resource-efficient agent designs.111 Data and codes released at https://github.com/qiancheng0/Open-SMARTAgent
[…]
footnote
@@
Cheng Qian1thanks:   indicates equal contribution., Emre Can Acikgoz1∗, Hongru Wang1, Xiusi Chen1, Avirup Sil2, Dilek Hakkani-Tür1, Gokhan Tur1, Heng Ji1 1University of Illinois Urbana-Champaign, 2IBM Research AI {chengq9, acikgoz2, hengji}@illinois.edu
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) (Ouyang et al., 2022; Team et al., 2023; Dubey et al., 2024) have led to remarkable improvements in reasoning capabilities, driving progress in diverse domains such as coherent text composition (Wei et al., 2022a), code generation (Gao et al., 2023; Wang et al., 2025; Pan et al., 2024), complex logical deduction (Yao et al., 2023, 2024), and nuanced natural language understanding (Wang et al., 2023; Yu et al., 2024; Wu et al., 2025). However, challenges remain, such as the inability to handle real-time information Yu and Ji (2024), provide accurate mathematical results (Lu et al., 2022), and fully comprehend human intentions Qian et al. (2024b). These limitations highlight the need for LLMs to leverage external tools (Schick et al., 2023; Qin et al., 2023; Yuan et al., 2024; Qian et al., 2024a), enabling them to function as agents capable of assisting users in diverse tasks (Qin et al., 2024; Xi et al., 2023). Effective tool use and reasoning are thus complementary, each enhancing the other to overcome current shortcomings.
@@
Figure 1: An illustration of human metacognition in problem-solving. The user recalls Tim Cook’s association with Apple based on prior knowledge, as it is a slow-changing fact. However, recognizing the newest chip as a fast-changing fact beyond knowledge scope, they switch to online search for up-to-date information.
@@ 2 Related Work
Recent studies highlight that while LMs excel at standard tasks, they struggle to recognize and acknowledge the limits of their knowledge Yin et al. (2023); Kadavath et al. (2022). To address this gap, the concept of knowledge boundary has been introduced to define the limits of knowledge in LLMs Li et al. (2024); Amayuelas et al. (2023). Building on this, some research evaluates LMs’ self-awareness of their knowledge boundary through verbal probing (Kadavath et al., 2022) and fine-grained benchmarks (Yin et al., 2024), enabling LMs to determine whether a question is answerable. Other work focuses on mitigating hallucinations arising from the model’s unawareness of its limits through data augmentation Chen et al. (2023, 2024b), retrieval augmentation Ren et al. (2023), and confidence calibration (Xue et al., 2024). Additionally, Chen et al. (2024a) and Zhang et al. (2024) trained LLMs to express their knowledge boundaries, enabling them to answer known questions and admit ignorance for unknown ones. Our work aligns with these studies but focuses on enhancing agents’ awareness of their knowledge boundaries to enable wiser tool use and more efficient task handling.
@@
Integrating tool use into LLMs has gained significant attention as a way to complement parametric knowledge and enhance decision-making Qin et al. (2023); Qu et al. (2025). Some research focuses on enabling LLMs to access external tools to overcome knowledge limitations (Qin et al., 2024; Qian et al., 2024d), including up-to-date information (Vu et al., 2023; Wang et al., 2024b) and domain-specific expertise (Ling et al., 2023; Wang et al., 2024a). Others explore tool creation (Qian et al., 2023; Cai et al., 2024) and external module integration (Qian et al., 2024c) to improve tool learning robustness. Despite these, a key challenge lies in evaluating and enhancing LLMs’ ability to determine when and which tools to use. Benchmarks like MetaTool Huang et al. (2023) and WTU-EVAL Ning et al. (2024) highlight LLMs’ struggles with unnecessary or incorrect tool usage, while dynamic frameworks Wang et al. (2024c); Shen et al. (2024) propose adaptively invoking tools based on internal uncertainty thresholds. Unlike prior works, SMART rigorously defines and measures tool overuse, addressing overreliance on tools despite intrinsic reasoning capabilities. We optimize the balance between parametric knowledge and tool use, reducing overuse while enhancing performance.
@@ 3 Preliminaries
Definition of Tool Overuse.
Tool overuse refers to the excessive reliance on external tools when an agent model could have successfully completed the task using its parametric knowledge alone. Formally, let $Q$ be the total set of questions, and let $P$ be the subset of questions that the model can correctly answer without using any tools. The model’s intrinsic reasoning capability is then given by $\alpha=\frac{|P|}{|Q|}$. Now, suppose that when provided with access to tools, the model chooses to invoke at least one tool on a fraction $\beta$ of these questions in $P$. The Tool Overuse Rate is then defined as:
@@
Figure 3: Example cases respectively demonstrating tool overuse in XAgent and AgentGPT.
Experiments on LLMs.
We first experiment with Llama-3.1-8B (Dubey et al., 2024) and Mistral-7B (Jiang et al., 2023) on the GSM8K test set (Cobbe et al., 2021). Each test question is presented under two conditions: i) the model reasons through the question normally and provides a final answer without using tools, and ii) the model has access to tools and independently decides whether to use them (see Section A.2). The statistics in Figure 2 reveal two key insights. First, both models exhibit significant tool overuse, with Llama’s rate exceeding 50%. Second, in some cases, tool use leads to incorrect answers, even for questions the model could have solved correctly without external assistance. This highlights how excessive reliance on tools can introduce unnecessary complexity and degrade performance.
Experiments on LM-driven Agents.
In addition to LLMs, we also experiment with two agent systems: XAgent (XAgent-Team, 2023) and AgentGPT (Team, 2024), both designed for complex problem-solving and driven by closed-source GPT models. We sampled 50 queries from the GSM8K test set that can be answered correctly without tools (see Section A.1) and instructed the models to use tools only when necessary. The results in Table 1 show that, despite being equipped with various tools, both agent systems still tend to use them unnecessarily, significantly slowing down problem-solving (about 10x slower than using GPT alone). In the case study shown in Figure 3, we identify issues like XAgent redundantly saving results and AgentGPT unnecessarily invoking a code-writing tool after generating answers. These observations underscore the need to address our core research question: How can we calibrate agent models to balance tool use and parametric reasoning, mitigating tool overuse while preserving utility?
@@ 5.1 Settings
For reasoning without tools, the model generates a response including the final answer. For tool-reliant reasoning, the inference is performed in an interactive manner: in each round, if a tool call is detected, we parse and execute it, integrating the tool’s output and reasoning into the input. This repeats until the final answer is reached. See Appendix C for details.
@@ 5.2 Main Results
SMARTAgent solves tasks efficiently. Compared to the base model in Table 3, which autonomously decides whether to use tools, SMARTAgent reduces tool usage time per query by 24% on average. At the same time, its performance improves by over 37% across models compared to the best baseline. This demonstrates SMARTAgent’s efficiency in tool use, achieving higher results while relying less on external resources.
7B-scale SMARTAgent can outperform GPT-4o baselines. Despite being much smaller, the 7B- and 8B-scale SMARTAgent models can outperform GPT-4o and its 70B counterpart in Time and Intention domains while using fewer tool calls, showcasing their efficient tool use. In Math, where reasoning scales with model size, SMARTAgent lags behind larger models but remains competitive against baselines using the same architecture. These results demonstrate that strategic tool use can bridge the gap between model size and performance, making SMARTAgent a resource-efficient yet powerful alternative.
SMARTAgent generalizes to OOD settings. As shown in Table 4, SMARTAgent effectively reduces tool calls while achieving better overall performance on OOD test benchmarks. Notably, SMARTAgent makes only one-fifth the number of tool calls compared to the base model in MINTQA, where tool prompting often leads to excessive reliance and decreased accuracy.
Improper tool uses degrade performance. In the MINTQA and Math domain data, we find that arbitrary tool use can degrade performance compared to standard chain-of-thought reasoning. This aligns with our argument in Figure 3 that excessive tool reliance can introduce unpredictable side effects, causing models to struggle with interactive tool calls. As a result, inference may become prolonged over multiple rounds, ultimately leading to incorrect answers. Additionally, we observe that larger-scale models, including GPT-4o, use tools less frequently in the Intention domain data, resulting in a greater performance drop than even the 7B-scale SMARTAgent. This may stem from their overconfidence in assisting users, leading them to overlook specific user preferences.
SMARTAgent achieves near-optimal tool use. Datasets such as Time and MINTQA contain up-to-date knowledge necessitating tool use. Ideally, at least one tool call per query is required for a correct answer, and SMARTAgent consistently maintains an average close to one, reflecting near-optimal efficiency. Similarly, in the Intention domain, where queries contain two to four missing details, SMARTAgent invokes tools three times per query, aligning with the expected need.
Table 5: Error analysis of common task failure causes, highlighting several error aspects: timing of tool calls, hyper-parameters provided, and handling of tool call feedback, with explanations and examples.
Table 6: Statistics of tool overuse on GSM8K and Math domain test data, following the definition in Section 3.
Figure 6: A case study comparing Tool Prompting and SMARTAgent reveals that SMARTAgent excels by creating thorough plans, demonstrating awareness of its limitations, and making justified, wiser tool call decisions.
@@ 5.3 Analysis and Case Studies
SMARTAgent effectively reduces tool overuse. Beyond measuring tool use per query, we calculate the tool overuse rate, as defined in Section 3, and report results in Table 6 for GSM8K and Math domain test data. Notably, SMARTAgent reduces unnecessary tool calls by up to 50% compared to prompting the base model with tool access. However, despite this reduction, tool overuse persists, which we further examine in error analysis.
@@
Error analysis. We provide error analysis in Table 5, highlighting common failure causes. Tool prompting leads to errors across all categories, while SMARTAgent reduces repetitive calls and improves argument accuracy. However, feedback neglect still causes tool invocation failures, particularly with the Code tool, and excessive caution in ensuring calculation accuracy adds overhead. This mirrors human task-solving, where we sometimes rely on calculators despite knowing the steps. Future work may explore balancing convenience, budget, and efficiency to enhance decision-making.
@@
Case Study. In Figure 6, we compare the solution chains of SMARTAgent and the base model with tool prompting. SMARTAgent demonstrates logical planning, context corroboration, and an awareness of its limitations and knowledge boundaries, with clear justifications for its decisions. This metacognitive approach closely mirrors human reasoning processes, making SMARTAgent’s reasoning more interpretable and significantly reducing tool use overhead.
Confidence Validation Experiment. To assess how effectively SMARTAgents choose between internal reasoning and tool invocation, we conducted experiments by injecting special tokens and analyzing the model’s confidence in selecting each one.
We trained the model on the Time and Intention domains, introducing three special tokens: “[[Reasoning]]” for internal reasoning, “[[AskUser]]” for the AskUser tool, and “[[Search]]” for the Search tool. These tokens were prepended at the start of each step during training to guide the model’s decision-making. See Section C.5 for details.
For evaluation, we sampled 50 decision steps from the test split of both domains and examined the logits of the selected tokens to measure confidence. Decisions were categorized as correct or wrong based on whether the model’s choice aligned with the ground truth. As shown in Figure 7, the model consistently exhibited higher confidence in correct decisions than incorrect ones. This demonstrates that SMART effectively boosts confidence in accurate decision-making, reinforcing its ability to distinguish when to rely on internal knowledge or external tools.
6 Discussions and Future Work
@@
Agent’s improper tool usage. Our empirical analysis reveals a notable phenomenon of tool overuse, where agents frequently rely on external tools even when internal knowledge is sufficient. This over-reliance likely arises from two factors: i) the agent’s uncertainty about its own capabilities, and ii) the perceived ease of external lookups compared to internal reasoning. We also observe instances of tool underuse, especially in large-scale models like GPT-4o and Llama-70B, where agents neglect to call essential tools, possibly due to misjudging the complexity of the task. Both overuse and underuse contribute to concerns over computational efficiency and solution accuracy. Future research could explore methods to better balance these trade-offs, such as by introducing explicit resource constraints or budgets for tool calls.
@@
Mechanisms behind human and LM’s decision-making. Cognitive science suggests that human decision-making arises from both intuitive judgments and reflective strategies. Similarly, in LMs, problem-solving is influenced by implicit heuristics (e.g., memorized patterns) and explicit tool-using behaviors. When tools are available, LMs often default to external queries, akin to humans seeking external confirmation when uncertain. However, LMs lack self-monitoring and rely on external or data-driven cues to determine when to trust their internal knowledge. Developing frameworks that integrate implicit heuristics with explicit reasoning could lead to more adaptive and efficient LM decision-making.
@@
Enhancement of model’s self-awareness. Our data-driven calibration strategy, which provides explicit rationales for when to rely on internal knowledge versus external tools, shows promising results. Other approaches, such as confidence probing via logits, integration of specialized self-checking modules, or reinforcement learning from feedback, might also refine tool usage thresholds. Future research could investigate how these signals affect the model’s internal distributions and identify representations that capture the awareness of boundaries. Additionally, iterative or in-context learning could allow real-time metacognitive calibration, offering a more efficient safeguard against both overuse and underuse of resources.

### v2

@@ Abstract
Current Large Language Model (LLM) agents demonstrate strong reasoning and tool use capabilities, but often lack self-awareness, failing to balance these approaches effectively. This imbalance leads to Tool Overuse, where models unnecessarily rely on external tools for tasks solvable with parametric knowledge, increasing computational overhead. Inspired by human metacognition, we introduce SMART (Strategic Model-Aware Reasoning with Tools), a paradigm that enhances an agent’s self-awareness to optimize task handling and reduce tool overuse. To support this paradigm, we introduce SMART-ER, a dataset spanning three domains, where reasoning alternates between parametric knowledge and tool-dependent steps, with each step enriched by rationales explaining when tools are necessary. Through supervised training, we develop SMARTAgent, a family of models that dynamically balance parametric knowledge and tool use. Evaluations show that SMARTAgent reduces tool use by 24% while improving performance by over 37%, enabling 7B-scale models to match its 70B counterpart and GPT-4o. Additionally, SMARTAgent generalizes to out-of-distribution test data like GSM8K and MINTQA, maintaining accuracy with just one-fifth the tool calls. These highlight the potential of strategic tool use to enhance reasoning, mitigate overuse, and bridge the gap between model size and performance, advancing intelligent and resource-efficient agent designs. All the data and codes are released111 https://github.com/qiancheng0/Open-SMARTAgent.
@@
Cheng Qian1thanks:  Indicates equal contribution., Emre Can Acikgoz1∗, Hongru Wang1thanks:  Mentorship, Xiusi Chen1, Avirup Sil2, Dilek Hakkani-Tür1, Gokhan Tur1, Heng Ji122footnotemark: 2 1University of Illinois Urbana-Champaign, 2IBM Research AI {chengq9, acikgoz2, hengji}@illinois.edu
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) (Ouyang et al., 2022; Team et al., 2023; Dubey et al., 2024) have led to remarkable improvements in reasoning capabilities, driving progress in diverse domains such as coherent text composition (Wei et al., 2022a), code generation (Gao et al., 2023; Wang et al., 2025b; Pan et al., 2024), complex logical deduction (Yao et al., 2023, 2024), and nuanced natural language understanding (Wang et al., 2023; Yu et al., 2024; Wu et al., 2025). However, challenges remain, such as the inability to handle real-time information Yu and Ji (2024), model real-world challenges (Qian et al., 2025b), provide accurate mathematical results (Lu et al., 2022), and fully comprehend human intentions Qian et al. (2024b). These limitations highlight the need for LLMs to leverage external tools (Schick et al., 2023; Qin et al., 2023; Yuan et al., 2024; Qian et al., 2024a), enabling them to function as agents capable of assisting users in diverse tasks (Qin et al., 2024; Xi et al., 2023). Effective tool use and reasoning are thus complementary, each enhancing the other to overcome current shortcomings.
@@
Figure 1: An illustration of human metacognition: The user recalls Tim Cook’s role from prior knowledge (a slow-changing fact), but uses online search to find the latest chip info (a fast-changing fact).
@@ 2 Related Work
Recent studies highlight that while LMs excel at standard tasks, they struggle to recognize and acknowledge the limits of their knowledge Yin et al. (2023); Qian et al. (2023b); Kadavath et al. (2022). To address this gap, the concept of knowledge boundary has been introduced to define the limits of knowledge in LLMs Li et al. (2024); Amayuelas et al. (2023). Building on this, some research evaluates LMs’ self-awareness of their knowledge boundary through verbal probing (Kadavath et al., 2022) and fine-grained benchmarks (Yin et al., 2024), enabling LMs to determine whether a question is answerable. Other work focuses on mitigating hallucinations arising from the model’s unawareness of its limits through data augmentation Chen et al. (2023, 2024b), retrieval augmentation Ren et al. (2023), and confidence calibration (Xue et al., 2024). Additionally, Chen et al. (2024a) and Zhang et al. (2024) trained LLMs to express their knowledge boundaries, enabling them to answer known questions and admit ignorance for unknown ones. Recently, reinforcement learning has been increasingly explored as a means to help models recognize their knowledge boundaries to guide more efficient decisions (Qian et al., 2025a; Wang et al., 2025a). Our work aligns with these studies and focuses on enhancing agents’ awareness for wiser tool use.
@@
Integrating tool use into LLMs has gained significant attention as a way to complement parametric knowledge and enhance decision-making Qin et al. (2023); Qu et al. (2025). Some research focuses on enabling LLMs to access external tools to overcome knowledge limitations (Qin et al., 2024; Qian et al., 2024d), including up-to-date information (Vu et al., 2023; Wang et al., 2024b) and domain-specific expertise (Ling et al., 2023; Wang et al., 2024a). Others explore tool creation (Qian et al., 2023a; Cai et al., 2024) and external module integration (Qian et al., 2024c) to improve tool learning robustness. Despite these, a key challenge lies in evaluating and enhancing LLMs’ ability to determine when and which tools to use. Benchmarks like MetaTool Huang et al. (2023) and WTU-EVAL Ning et al. (2024) highlight LLMs’ struggles with unnecessary or incorrect tool usage, while dynamic frameworks Wang et al. (2024c); Shen et al. (2024) propose adaptively invoking tools based on internal uncertainty thresholds. Unlike prior works, SMART rigorously defines tool overuse and addresses it by optimizing the balance of internal knowledge and tool use.
@@ 3 Preliminaries
Definition of Tool Overuse. Tool overuse refers to the excessive reliance on external tools when an agent model could have successfully completed the task using its parametric knowledge alone. Formally, let $Q$ be the total set of questions, and let $P$ be the subset of questions that the model can correctly answer without using any tools. The model’s intrinsic reasoning capability is then given by $\alpha=\frac{|P|}{|Q|}$. Now, suppose that when provided with access to tools, the model chooses to invoke at least one tool on a fraction $\beta$ of these questions in $P$. The Tool Overuse Rate is then defined as:
@@
Experiments on LLMs. We first experiment with Llama-3.1-8B (Dubey et al., 2024) and Mistral-7B (Jiang et al., 2023) on the GSM8K test set (Cobbe et al., 2021). Each test question is presented under two conditions: i) the model reasons through the question normally and provides a final answer without using tools, and ii) the model has access to tools and independently decides whether to use them (see Section A.2). The statistics in Figure 2 reveal two key insights. First, both models exhibit significant tool overuse, with Llama’s rate exceeding 50%. Second, in some cases, tool use leads to incorrect answers, even for questions the model could have solved correctly without external assistance. This highlights how excessive reliance on tools can introduce unnecessary complexity and degrade performance.
Figure 3: Example cases on XAgent and AgentGPT’s tool overuse.
Experiments on LM-driven Agents. In addition to LLMs, we also experiment with two agent systems: XAgent (Team, 2023) and AgentGPT (Team, 2024), both designed for complex problem-solving and driven by closed-source GPT models. We sampled 50 queries from the GSM8K test set that can be answered correctly without tools (see Section A.1) and instructed the models to use tools only when necessary. The results in Table 1 show that, despite being equipped with various tools, both agent systems still tend to use them unnecessarily, significantly slowing down problem-solving (about 10x slower than using GPT alone). We further provide a case study in Section A.1 highlighting issues such as XAgent redundantly saving results to files and AgentGPT unnecessarily invoking a code-writing tool after generating an answer. These observations underscore the need to address our core research question: How can we calibrate agent models to balance tool use and parametric reasoning, mitigating tool overuse while preserving utility?
@@ 5.1 Settings
For reasoning without tools, the model generates a response including the final answer. For tool-reliant reasoning, the inference is interactive: in each round, if a tool call is detected, we parse and execute it, integrating the tool’s output and reasoning into the input. This repeats until the final answer is reached. See Appendix C for details.
@@ 5.2 Main Results
SMARTAgent solves tasks efficiently.
Compared to the base model in Table 3, which autonomously decides whether to use tools, SMARTAgent reduces tool usage time per query by 24% on average. At the same time, its performance improves by over 37% across models compared to the best baseline. This demonstrates SMARTAgent’s efficiency in tool use, achieving higher results while relying less on external resources.
7B-scale SMARTAgent can outperform GPT-4o baselines.
Despite being much smaller, the 7B- and 8B-scale SMARTAgent models can outperform GPT-4o and its 70B counterpart in Time and Intention domains while using fewer tool calls, showcasing their efficient tool use. In Math, where reasoning scales with model size, SMARTAgent lags behind larger models but remains competitive against baselines using the same architecture. These results demonstrate that strategic tool use can bridge the gap between model size and performance, making SMARTAgent a resource-efficient yet powerful alternative.
SMARTAgent generalizes to OOD settings.
As shown in Table 4, SMARTAgent effectively reduces tool calls while achieving better overall performance on OOD test benchmarks. Notably, SMARTAgent makes only one-fifth the number of tool calls compared to the base model in MINTQA, where tool prompting often leads to excessive reliance and decreased accuracy.
Improper tool uses degrade performance.
In the MINTQA and Math domain data, we find that arbitrary tool use can degrade performance compared to standard chain-of-thought reasoning. This aligns with our argument in Table 1 that excessive tool reliance can introduce unpredictable side effects, causing models to struggle with interactive tool calls. As a result, inference may become prolonged over multiple rounds, ultimately leading to incorrect answers. Additionally, we observe that larger-scale models, including GPT-4o, use tools less frequently in the Intention domain data, resulting in a greater performance drop than even the 7B-scale SMARTAgent. This may stem from their overconfidence in assisting users, leading them to overlook specific user preferences.
SMARTAgent achieves near-optimal tool use.
Datasets such as Time and MINTQA contain up-to-date knowledge necessitating tool use. Ideally, at least one tool call per query is required for a correct answer, and SMARTAgent consistently maintains an average close to one, reflecting near-optimal efficiency. Similarly, in the Intention domain, where queries contain two to four missing details, SMARTAgent invokes tools three times per query, aligning with the expected need.
Table 5: Error analysis of common task failure causes, with explanations and examples.
Table 6: Statistics on tool overuse, defined in Figure 2.
@@
Figure 7: Case study comparing the performance of Tool Prompting and SMARTAgent.
@@ 5.3 Analysis and Case Studies
Beyond measuring tool use per query, we calculate the tool overuse rate, as defined in Figure 2, and report results in Table 6 for GSM8K and Math domain test data. Notably, SMARTAgent reduces unnecessary tool calls by up to 50% compared to prompting the base model with tool access. However, despite this reduction, tool overuse persists, which we further examine in error analysis.
Error analysis.
@@
We provide error analysis in Table 5, highlighting common failure causes. Tool prompting leads to errors across all categories, while SMARTAgent reduces repetitive calls and improves argument accuracy. However, feedback neglect still causes tool invocation failures, particularly with the Code tool, and excessive caution in ensuring calculation accuracy adds overhead. This mirrors human task-solving, where we sometimes rely on calculators despite knowing the steps. Future work may explore balancing convenience, budget, and efficiency to enhance decision-making.
Case Study.
@@
In Figure 7, we compare the solution chains of SMARTAgent and the base model with tool prompting. SMARTAgent demonstrates logical planning, context corroboration, and an awareness of its limitations and knowledge boundaries, with clear justifications for its decisions. This metacognitive approach closely mirrors human reasoning processes, making SMARTAgent’s reasoning more interpretable and significantly reducing tool use overhead.
Confidence Validation Experiment.
To evaluate SMARTAgents’ ability to choose between internal reasoning and tool invocation, we conducted experiments using special tokens to analyze decision confidence. Specifically, we trained the model on Time and Intention domains, introducing special tokens: “[[Reasoning]]” for internal reasoning, “[[AskUser]]” for the AskUser tool, and “[[Search]]” for the Search tool. These tokens, prepended at each step, guided decision-making during training (see Section C.5). For evaluation, we sampled 50 decision steps from both domains’ test splits, measuring confidence via token logits. Decisions were categorized as correct or incorrect based on alignment with ground truth. As shown in Figure 6, the model exhibited higher confidence in correct decisions, demonstrating SMART’s effectiveness in boosting confidence and distinguishing between internal knowledge and tool use.
6 Discussions
Agent’s improper tool usage.
@@ 6 Discussions
Our empirical analysis reveals a notable phenomenon of tool overuse, where agents frequently rely on external tools even when internal knowledge is sufficient. This over-reliance likely arises from two factors: i) the agent’s uncertainty about its own capabilities, and ii) the perceived ease of external lookups compared to internal reasoning. We also observe instances of tool underuse, especially in large-scale models like GPT-4o and Llama-70B, where agents neglect to call essential tools, possibly due to misjudging the complexity of the task. Both overuse and underuse contribute to concerns over computational efficiency and solution accuracy. Future research could explore methods to better balance these trade-offs, such as by introducing explicit resource constraints or budgets for tool calls.
Mechanisms behind human and LM’s decision-making.
@@
Cognitive science suggests that human decision-making arises from both intuitive judgments and reflective strategies. Similarly, in language models (LMs), problem-solving is influenced by implicit heuristics (e.g., memorized patterns) and explicit tool-using behaviors. When tools are available, LMs often default to external queries, akin to humans seeking external confirmation when uncertain. However, unlike humans, LMs lack self-monitoring and rely on external or data-driven cues to determine when to trust their internal knowledge. Developing frameworks that integrate implicit heuristics with explicit reasoning could lead to more adaptive and efficient decision-making in LMs.
Enhancement of model’s self-awareness.
@@
Our data-driven calibration strategy, which provides explicit rationales for when to rely on internal knowledge versus external tools, shows promising results. Other approaches, such as confidence probing via logits, integration of specialized self-checking modules, or reinforcement learning from feedback, might also refine tool usage thresholds. Future research could investigate how these signals affect the model’s internal distributions and identify representations that capture the awareness of boundaries. Additionally, iterative or in-context learning could allow real-time metacognitive calibration, offering a more efficient safeguard against both overuse and underuse of resources.